# Learning in Chaos: Efficient Autoscaling and Self-healing for Distributed Training at the Edge

### v1

@@ VI-D Scale-in and Connect / Disconnect-link Delays of Chaos
In the above, we focused on scale-out because it requires pulling training state from other nodes, leading to noticeable communication delays. In contrast, other primitives, scale-in, connect-link, and disconnect-link, only exchange tiny control messages, and they overlap with costly operations like all-reduce and gradient computation, so their delays are hidden. In this experiment, we keep the number of nodes fixed and use connect-link to add a link between two random nodes. The results are shown in Fig. 9. The connect-link delay remains under 1 millisecond and is unaffected by cluster size. This result also applies to scale-in and disconnect-link. As a result, there is little value in comparing them across systems since they are never a bottleneck and all perform well in practice.
VI-E How Scale-in/out Affects Convergence?
@@
Adding or removing nodes brings in or takes away training data, so in this experiment, we explore how scale-out and scale-in affect model convergence, including ResNet101, AlexNet, and VGG11. Their accuracy curves are shown in Figs. 11 and 12, respectively. For scale-out, training starts with 4 nodes, and a new node is added at the 100th epoch (green). Before the new node joins, the scale-out curve (green) closely tracks the 4-node curve (blue), and it smoothly shifts toward the 5-node curve (orange) after the new node joins. For scale-in, training starts with 5 nodes, and one node is removed at the 100th epoch (green). Before a node leaves, the scale-in curve follows the 5-node curve and gradually shifts toward the 4-node curve afterward. Even after the node leaves, the model retains knowledge from its data. The same pattern was found when fine-tuning GPT-2 with LoRA, as shown in Figs. 13 and 14, but nodes are added or removed at epoch 3. This shows Chaos also works well for LLM fine-tuning.
Notably, both scale-out and scale-in curves stay smooth during node changes, with no sharp fluctuations, because deep models are robust to small changes in training data, minor additions or removals do not disrupt convergence. And since we change only one node at a time, the noise is further diluted by global averaging. This feature makes Chaos well-suited for edge DML training and fine-tuning, where nodes are free to join and leave, usually one by one, following a Poisson process, instead of in batches like on the cloud.
VI-F Ablation Study on State Replication Mechanism
State replication is key to a fast scale-out. In this experiment, we conduct ablation studies on the state replication mechanism of EDL+, Autoscaling, and Chaos to evaluate how effective our multi-neighbor replication is versus single-source and multi-source replication, and how small the optimality gap is between our greedy shard assignment and the optimum.
The multi-neighbor approach is inherently superior. In this study, the single-source approach picks the fastest neighbor, while multi-source and our multi-neighbor approaches use equal shard splits across nodes. As shown in Fig. 15, the multi-neighbor approach offers much lower scale-out delays than both single- and multi-source approaches, and in most cases, multi-source is even slower than single-source. This matches the analysis in Fig. 1. As discussed in Section II-B, relying on a single source creates bandwidth and load bottlenecks that only get worse as models and clusters scale up. While with multi-source replication, it still suffers from rapidly increasing delays due to redundant traffic and multi-hop forwarding. Multi-neighbor replication improves it by pulling only from neighbors, cutting out redundant traffic and long network paths, making it the standout approach.
Our greedy assignment is empirically near-optimal. By default, multi-neighbor replication assigns equal shard sizes to all neighbors, which we treat as the worst-case upper bound. We use brute-force search to find the global optimum as a lower bound. The closer we are to this lower bound, the better the shard assignment. As shown in Fig. 16, our greedy assignment stays within 0.5%–29% of the optimal, matching the analysis in Section III-B. This confirms that Algorithm 2 is easy-to-use, fast, and effective, making it a top choice for real-world deployment.
By combining the multi-neighbor replication with a smart shard scheduler, Chaos keeps scale-out delays lower and more stable, and can be even faster as the cluster grows.
@@ VI-F Ablation Study on State Replication Mechanism
Figure 15: Comparison of scale-out delays for single-source, multi-source and multi-neighbor replication.
Figure 16: Comparison of equal, greedy, and optimal assignment for multi-neighbor replication.
@@ VII Conclusion
This paper presents Chaos, a resilient and scalable edge DML training system with built-in self-healing and autoscaling. It implements four primitives: scale-out, scale-in, connect-link, and disconnect-link, to handle node and link joins, exits, and failures. To speed up scale-out, Chaos features multi-neighbor replication with fast shard scheduling to reduce state replication delay. A built-in cluster monitor tracks resource and topology changes, feeding real-time network data to the scheduler for scaling decisions. Through peer negotiation protocols, Chaos handles scaling events in a self-governed, decentralized manner, but not centrally managed by an admin. Experiments show that Chaos consistently achieves much lower scale-out delays than Pollux, EDL, and Autoscaling, and handles scale-in, connect-link, and disconnect-link events in under 1 millisecond, making it smoother to handle joins, exits, and failures. It also delivers the lowest idle time, showing superior resource use and scalability as the cluster grows.

### v2

@@ VI-D Scale-in, Connect- and Disconnect-link Delays in Chaos
In the above, we focus on scale-out because it requires state replication and has noticeable communication delays. However, other scaling primitives, including scale-in, connect-link, and disconnect-link, exchange only tiny control messages, and their delay overlaps with heavy operations like all-reduce and gradient computation, making them negligible. In this experiment, we add a link between two randomly chosen nodes. As shown in Fig. 11, the connect-link delay stays below 20ms, so it’s unnecessary to compare with baselines since it cannot become a bottleneck. This result also applies to scale-in and disconnect-link, because they have a similar traffic model.
(a) Scale-out
(b) Scale-in
Figure 14: Effects of scale-out, scale-in on GPT-2 convergence.
VI-E Ablation on How Scale-in and Out Affect Convergence?
@@
Adding or removing nodes changes the training data, so this experiment assesses how scale-out and scale-in affect model convergence. The accuracy curves for ResNet101, AlexNet, and VGG11 are shown in Figs. 12 and 13, respectively. For scale-out, training starts with 4 nodes, and a new node is added at the 100th epoch (green). Before the new node joins, the scale-out curve (green) closely tracks the 4-node curve (blue), and it smoothly shifts toward the 5-node curve (orange) after the new node joins. For scale-in, training starts with 5 nodes, and one node is removed at the 100th epoch (green). Before a node leaves, the scale-in curve follows the 5-node curve and gradually shifts toward the 4-node curve afterward. Even after the node leaves, the model retains knowledge from its data. The same pattern was found when fine-tuning GPT-2 with LoRA, as shown in Fig. 14. This shows that Chaos also works well for LLM fine-tuning. Moreover, both scale-out and scale-in curves remain smooth during node churn, without sharp fluctuations, because deep models are robust to small changes in training data, and minor additions or removals do not disrupt convergence. Since we change only one node at a time, the noise is further diluted by global averaging.
Fig. 15 shows the global accuracy curves on ResNet101. Training starts with 4 nodes and undergoes four churn events: one node joins at epoch 50; one node exits at epoch 97; one node joins at epoch 145; and one more node joins at epoch 194, ending with 6 nodes. The sub-figures plot accuracy gains over the 4-node baseline. After each join, accuracy steps up by a few points, reflecting the benefit of more parties (data). The node’s leave causes a small dip, but the model retains knowledge learned from the exited node, as the overall gain remains positive. These align with the local trend in earlier experiments. From a global view, training under churn (blue) converges to nearly the same accuracy as the 6-node run (green). This suggests that multi-party training under churn is feasible. We focus on exploring the effect of node churn on model accuracy, but methods for preserving memory under churn are complementary and left for future work. Nevertheless, prior studies like [34] have already shown promising solutions to mitigate forgetting in such settings.
VI-F Ablation on State Replication Mechanism
State replication is key to a fast scale-out. In this experiment, we run an ablation study to evaluate our replication mechanism, highlighting the benefits of multi-neighbor replication and our greedy sharding algorithm. The single-source method fetches from the fastest neighbor, whereas the multi-source method fetches from multiple nodes (not neighbors) with the model evenly sharded across them. In addition to our multi-neighbor replication with greedy sharding, we also provide a variant with equal sharding to observe the individual benefits of multi-neighbor replication and greedy sharding.
As shown in Fig. 16, multi-neighbor methods achieve lower scale-out delays than single- and multi-source methods, with multi-source often even slower than single-source. This aligns with Fig. 2: single-source suffers from bandwidth and load bottlenecks that worsen as models and clusters grow, while multi-source adds redundant traffic and multi-hop forwarding, leading to rapidly increasing delays. In contrast, multi-neighbor replication fetches only from neighbors, eliminating redundancy and long paths. With greedy sharding, Chaos keeps scale-out delays lower and more stable, and even faster as the cluster grows, by precisely balancing transmission loads across neighbors. Its sharding and assignment solution has been proven optimal in Section III-B.
Figure 15: Global accuracy curves on ResNet101. Green: fixed 6 nodes. Orange: fixed 4 nodes. Blue: 4-6 nodes with churn.
@@ VI-F Ablation on State Replication Mechanism
Figure 16: Comparison of single-source, multi-source, and multi-neighbor state replication.
@@ VII Conclusion
This paper proposes Chaos, a multi-party distributed training system with self-healing and autoscaling, enabling robust and elastic training under churn. It implements four scaling primitives: scale-out, scale-in, connect-link, and disconnect-link, to handle node and link joins, exits, and failures. To speed up autoscaling, we propose multi-neighbor replication with an optimal sharding and assignment mechanism to minimize scale-out delay. A built-in cluster monitor tracks resource and topology changes, feeding real-time network data to the scheduler for scaling decisions. Through peer negotiation protocols, Chaos handles scaling events in a fully self-governed manner. Experiments show that Chaos achieves substantially lower scale-out delay than Pollux, Elan, and Autoscaling, and handles scale-in, connect-link, and disconnect-link events within 20ms, making it smooth to handle the churn in a cross-region cluster. It also delivers the lowest idle time, showing superior resource use and scalability as the cluster grows.
Acknowledgments
We thank SAIBIT TECH for providing equipment support.